OpenStudy (anonymous):

Somebody please help me! Calculate the consumer's surplus at the indicated unit price (P bar) for the demand equation. (Round your answer to the nearest cent.) q= 20-0.05p^2; P bar=2

OpenStudy (anonymous):

P stand for the number 2. p^2 meaning you have to take the number p and multiply it 2 times. 20-0.05 = 19.95. (we have to multiply because when you see no sign between a problem like 20-0.05 p^2 that mean multiply.) 19.95*4= 79.8 so, q= 79.8

OpenStudy (anonymous):

the answer to this problem is ( q= 79.8 ).

OpenStudy (anonymous):

So once I get 79.8 for q do I then integrate?